I am having a problem with this word problem, I am not good at them and need some help.
Geometry: The length of a rectangle is 1 cm longer than its width. If the diagonial of the rectangle is 4cm what are the dimensions of the rectangle?
Thanks a million

Answer by scott8148(6628)

let w=width, so w+1=length ... the length, width, and diagonal form a right triangle with the diagonal as the hypotenuse

by Pythagoras, w^2+(w+1)^2=4^2 ... 2w^2+2w+1=16 ... 2w^2+2w-15=0 ...

so w=2.28 and w=-3.28 ... the negative value is not realistic ... length=3.28